# Equivalence of entropy solutions and gradient flows for pressureless 1D Euler systems

### v2

@@ 6.1. Discrete setup
Recalling the setting of Section 2, we have $n$ initially distinct particles located at positions $x_{i}$ with velocities $v_{i}$ and masses $m_{i}$. From their masses we can define a partition of $\Omega=(0,1)$ given by $10=\theta_{0}<\theta_{1}<\dots<\theta_{n-1}<\theta_{n}=$ 1 where
@@
This shows that $-\frac{\alpha}{2}(M(t,x)^{2}-M(t,x))$ can be written as $\int_{0}^{M(t,x)}A_{n}(m)\mathop{}\!\mathrm{d}m$ for the piecewise constant function
@@
At this stage, with the initial particle evolution defined, we still need to derive the corresponding conservation law (4.4) for this discrete setting. From now on, to emphasize that we are considering particle solutions, we will add the subscript $n$ to both the Lagrangian variables $X$ and $V$, the cumulative quantities $M$ and $Q$, and the primitive $\mathcal{V}^{0}$. Moreover, as the position $t\mapsto x_{i}(t)$ and velocity $t\mapsto v_{i}(t)$ of the $i$th particle depend on time, so will the corresponding $X_{n}(t,m)$, $V_{n}(t,m)$, $M_{n}(t,x)$ and $Q_{n}(t,x)$.
@@ 6.2. Dynamics prior to first collision
From the point of view of conservation laws, we see that the particles, through their representation by the piecewise constant mass distribution $M_{n}$ and the piecewise linear and continuous flux function $\mathcal{U}_{n}$, correspond exactly to Dafermos’ [18], or the front tracking [9, 25], method for scalar conservation laws, at least initially. This correspondence is exactly what was used to provide “sticky particle” solutions of the Euler system in [8]. Since $M_{n}(t,x)$ is piecewise constant, it will be a weak solution of (6.6), the discrete counterpart of the scalar conservation law (4.4), provided it satisfies the Rankine–Hugoniot condition (4.16) along the discontinuities $x_{i}(t)$. Moreover, it will be an entropy solution if it additionally satisfies the Oleĭnik E-condition (4.17) along the discontinuities, and in particular, it will be the unique $BV$ solution of this conservation law for $t\in[0,T]$. This is indeed the case, since along the $i$th trajectory we have a left and right state, $\theta_{i-1}$ and $\theta_{i}$ respectively, such that $[[M_{n}]]=\theta_{i}-\theta_{i-1}=m_{i}$ and $[[\mathcal{U}_{n}(t,M_{n})]]=m_{i}(v_{i}^{0}+ta_{i})=m_{i}u_{i}(t)$. Therefore, the Rankine–Hugoniot condition (4.16) is satisfied since
@@ 7.1.1. Computing the (non-)Lagrangian trajectories
$\tilde{X}(t,m)$ =$=\begin{cases}(1+t^{2})(m-\frac{1}{2})-t\operatorname*{sgn}(m-\frac{1}{2}),&\lvert m-\frac{1}{2}\rvert\geq\frac{t}{1+t^{2}},\\ 0,&\lvert m-\frac{1}{2}\rvert<\frac{t}{1+t^{2}}.\end{cases}$
@@
Combining the above, we see that at $t=1$, all mass is concentrated at the origin with zero velocity. For $10\leq t<$ 1, $\Omega_{X(t)}$ was contained in the set where $U$ is decreasing. For $t>1$ we have the opposite situation, and therefore
@@ 7.1.2. Entropy admissibility
where we have introduced the right-continuous sign function $\operatorname*{sgn}^{\!+}(x)$, i.e., $\operatorname*{sgn}(x)$ with the altered convention $\operatorname*{sgn}^{\!+}(0)=1$. Further computations, rather cumbersome in the case of $M$, show that the generalized inverses are piecewise smooth solutions of (4.4), with a jump discontinuity, or stationary shock, at the origin for $t>0$. Hence, to be weak and entropy solutions it is sufficient that they satisfy respectively the Rankine–Hugoniot condition and Oleĭnik E-condition at the shock. The symmetry of the flux function around $m=\frac{1}{2}$ yields $[[\mathcal{U}(t,\tilde{M}(t,0))]]=0=[[\mathcal{U}(t,M(t,0))]]$, so we see that both solutions satisfy the Rankine–Hugoniot condition at the stationary shock.
@@
Observe that $\mathcal{U}(t,m)$ is always concave on the set $\Omega_{X(t)}$, and since both this set and the flux function $\mathcal{U}(t,m)$ are symmetric around $m=\frac{1}{2}$, its lower convex envelope $\mathcal{U}_{\smallsmile}(t,m)$ on this set will be a horizontal chord connecting the endpoints. In consequence, we have $\mathcal{U}(t,m)>\mathcal{U}_{\smallsmile}(t,m)$ for $m\in\Omega_{X(t)}$, which in turn implies the Oleĭnik E-condition (4.17); compare this to the condition (7.2). On the other hand, on the set $\Omega_{\tilde{X}(t)}\setminus\Omega_{X(t)}$, which becomes nonempty only for $t>1$, $\mathcal{U}(t,m)$ is convex. Therefore, the lower convex envelope on the set $\Omega_{\tilde{X}(t)}$ will have a negative slope for $-\frac{t}{1+t^{2}}<m<-\frac{1}{2t}$ and a positive slope for $\frac{1}{2t}<m<\frac{t}{1+t^{2}}$, and this shows that the stationary shock of $\tilde{M}$ violates (4.17). These two solutions are compared in Figure 4, where we observe, as seen from the formulas, that the shock decreases faster for the entropy solution $M$.
@@ 7.1.3. Energy of trajectories
Hence, as the system tends to decrease its potential energy, for $t\geq 1$ there is no more dissipation of energy for the entropic solution $(X,V)$, as the kinetic and potential energy exactly balance one another. On the other hand, for $(\tilde{X},\tilde{V})$ the energy will increase again for $t>1$ and tend asymptotically to the initial value. Therefore, as in Example 6.1, we see that the Lagrangian, or entropy, solution provides a more physically desirable behavior in the sense that energy is nonincreasing.
@@
In the above we could also have considered a globally sticky solution where for $10\leq t<$ 1 the evolution is identical to $X$ and $\tilde{X}$, while for $t\geq 1$ all mass is concentrated in the stationary shock at the origin. This would also be a generalized Lagrangian solution, as the zero element is always contained in the tangent cone $T_{X}\mathscr{K}$, as well as a weak solution of the scalar conservation law since the Rankine–Hugoniot condition is satisfied for the generalized inverse. In addition, there would clearly be no more dissipation of energy. However, from analogous arguments as before, it fails to satisfy the entropy condition, as $\mathcal{U}(t,m)$ is no longer concave on the whole of $[0,1]$ for $t>1$.
@@ 7.2.4. Asymptotic behavior of the damped system
Let $(\rho,v)$ be the solution of (1.6) corresponding to the Lagrangian, or entropy, solution for initial data $\rho_{0}\in\mathscr{P}_{2}(\mathbb{R})$ and $v_{0}\in L^{2}(\mathbb{R},\rho_{0})$. Then the asymptotic solution is given by
@@
where we again have used (7.3). In the final expression, the second and third integral vanish due to Lemma 3.6 and (5.8); the second because of $U-V\in N_{X}\mathscr{K}\subset\mathscr{H}_{X}^{\perp}$, while the third vanishes for a.e. $t$ since $V(t)\in\mathscr{H}_{X(t)}$ for a.e. $t$. We note that $U-V\in N_{X}\mathscr{K}$ while $12m-$ 1 is nondecreasing, which allow us to combine the characterizations (3.6) and (3.9) to deduce that the final term is nonpositive, yielding
@@
Here $(X-\tilde{X},V-\tilde{V})$ corresponds to a uniform distribution centered at the center of mass for the system, and according to (7.7) this converges uniformly pointwise in $\Omega$ to the proposed steady state (7.8). From the fact that $-\tilde{X},\tilde{V}\in T_{X}\mathscr{K}$, (3.6), (5.8), $\tilde{U}-\tilde{V}=U-V\in N_{X}\mathscr{K}$ and $\int_{0}^{1}\tilde{X}\mathop{}\!\mathrm{d}m=0=\int_{0}^{1}\tilde{U}\mathop{}\!\mathrm{d}m$ we derive the following relations,
@@
The first identity in the second line holds for a.e. $t$ since $\tilde{V}(t)\in\mathscr{H}_{X(t)}$ for a.e. $t$, alternatively it holds for all times with an inequality since $\tilde{V}\in T_{X}\mathscr{K}$. From these relations, the evolution equations for $(\tilde{X},\tilde{V})$ and (7.6), we then, omitting the subscript $L^{2}(0,1)$ from norms and inner products for brevity, obtain

### v3

@@ 6.1. Discrete setup
Recalling the setting of Section 2, we have $n$ initially distinct particles located at positions $x_{i}$ with velocities $v_{i}$ and masses $m_{i}$. From their masses we can define a partition of $\Omega=(0,1)$ given by $0=\theta_{0}<\theta_{1}<\dots<\theta_{n-1}<\theta_{n}=1$ where
@@
This shows that $\frac{\alpha}{2}(M(t,x)^{2}-M(t,x))$ can be written as $\int_{0}^{M(t,x)}A_{n}(m)\mathop{}\!\mathrm{d}m$ for the piecewise constant function
@@
At this stage, with the initial particle evolution defined, we still need to derive the corresponding conservation law (4.4) for this discrete setting. From now on, to emphasize that we are considering particle solutions, we will add the subscript $n$ to both the Lagrangian variables $X$ and $V$, the distribution functions $M$ and $Q$, and the primitive $\mathcal{V}^{0}$. Moreover, as the position $t\mapsto x_{i}(t)$ and velocity $t\mapsto v_{i}(t)$ of the $i$th particle depend on time, so will the corresponding $X_{n}(t,m)$, $V_{n}(t,m)$, $M_{n}(t,x)$ and $Q_{n}(t,x)$.
@@ 6.2. Dynamics prior to first collision
From the point of view of conservation laws, we see that the particles, through their representation by the piecewise constant mass distribution $M_{n}$ and the piecewise linear and continuous flux function $\mathcal{U}_{n}$, correspond exactly to Dafermos’ [18], or the front tracking [9, 25], method for scalar conservation laws, at least initially. This correspondence is exactly what was used to provide “sticky particle” solutions of the Euler system in [8]. Since $M_{n}(t,x)$ is piecewise constant, it will be a weak solution of (6.6), the discrete counterpart of the scalar conservation law (4.4), provided it satisfies the Rankine–Hugoniot condition (4.16) along the discontinuities $x_{i}(t)$. Moreover, it will be an entropy solution if it additionally satisfies the Oleĭnik E-condition (4.17) along the discontinuities, and in particular, it will be the unique $BV$ solution of this conservation law for $t\in[0,T]$. This is indeed the case, since along the $i$th trajectory we have a left and right state, $\theta_{i-1}$ and $\theta_{i}$ respectively, such that $M_{n}]]=\theta_{i}-\theta_{i-1}=m_{i}$ and $\mathcal{U}_{n}(t,M_{n})]]=m_{i}(v_{i}^{0}+ta_{i})=m_{i}u_{i}(t)$. Therefore, the Rankine–Hugoniot condition (4.16) is satisfied since
@@ 7.1.1. Computing the (non-)Lagrangian trajectories
$\tilde{X}(t,m)$ $=\begin{cases}(1+t^{2})(m-\frac{1}{2})-t\operatorname*{sgn}(m-\frac{1}{2}),&\lvert m-\frac{1}{2}\rvert\geq\frac{t}{1+t^{2}},\\ 0,&\lvert m-\frac{1}{2}\rvert<\frac{t}{1+t^{2}}.\end{cases}$
@@
Combining the above, we see that at $t=1$, all mass is concentrated at the origin with zero velocity. For $0\leq t<1$, $\Omega_{X(t)}$ was contained in the set where $U$ is decreasing. For $t>1$ we have the opposite situation, and therefore
@@ 7.1.2. Entropy admissibility
where we have introduced the right-continuous sign function $\operatorname*{sgn}^{\!+}(x)$, i.e., $\operatorname*{sgn}(x)$ with the altered convention $\operatorname*{sgn}^{\!+}(0)=1$. Further computations, rather cumbersome in the case of $M$, show that the generalized inverses are piecewise smooth solutions of (4.4), with a jump discontinuity, or stationary shock, at the origin for $t>0$. Hence, to be weak and entropy solutions it is sufficient that they satisfy respectively the Rankine–Hugoniot condition and Oleĭnik E-condition at the shock. The symmetry of the flux function around $m=\frac{1}{2}$ yields $\mathcal{U}(t,\tilde{M}(t,0))]]=0=[[\mathcal{U}(t,M(t,0))]]$, so we see that both solutions satisfy the Rankine–Hugoniot condition at the stationary shock.
@@
Observe that $\mathcal{U}(t,m)$ is always concave on the set $\Omega_{X(t)}$, and since both this set and the flux function $\mathcal{U}(t,m)$ are symmetric around $m=\frac{1}{2}$, its lower convex envelope $\mathcal{U}_{\smallsmile}(t,m)$ on this set will be a horizontal chord connecting the endpoints. In consequence, we have $\mathcal{U}(t,m)>\mathcal{U}_{\smallsmile}(t,m)$ for $m\in\Omega_{X(t)}$, which in turn implies the Oleĭnik E-condition (4.17); compare this to the condition (7.2). On the other hand, on the set $\Omega_{\tilde{X}(t)}\setminus\Omega_{X(t)}$, which becomes nonempty only for $t>1$, $\mathcal{U}(t,m)$ is convex. Therefore, the lower convex envelope on the set $\Omega_{\tilde{X}(t)}$ will have a negative slope for $\frac{t}{1+t^{2}}<m<-\frac{1}{2t}$ and a positive slope for $\frac{1}{2t}<m<\frac{t}{1+t^{2}}$, and this shows that the stationary shock of $\tilde{M}$ violates (4.17). These two solutions are compared in Figure 4, where we observe, as seen from the formulas, that the shock decreases faster for the entropy solution $M$.
@@ 7.1.3. Energy of trajectories
[…]
@@
In the above we could also have considered a globally sticky solution where for $0\leq t<1$ the evolution is identical to $X$ and $\tilde{X}$, while for $t\geq 1$ all mass is concentrated in the stationary shock at the origin. This would also be a generalized Lagrangian solution, as the zero element is always contained in the tangent cone $T_{X}\mathscr{K}$, as well as a weak solution of the scalar conservation law since the Rankine–Hugoniot condition is satisfied for the generalized inverse. In addition, there would clearly be no more dissipation of energy. However, from analogous arguments as before, it fails to satisfy the entropy condition, as $\mathcal{U}(t,m)$ is no longer concave on the whole of $[0,1]$ for $t>1$.
@@ 7.2.4. Asymptotic behavior of the damped system
Let $(\rho,v)$ be the solution of (1.7) corresponding to the Lagrangian, or entropy, solution for initial data $(\rho_{0},v_{0})$ satisfying (1.2). Then the asymptotic solution is given by
@@
where we again have used (7.3). In the final expression, the second and third integral vanish due to Lemma 3.6 and (5.8); the second because of $U-V\in N_{X}\mathscr{K}\subset\mathscr{H}_{X}^{\perp}$, while the third vanishes for a.e. $t$ since $V(t)\in\mathscr{H}_{X(t)}$ for a.e. $t$. We note that $U-V\in N_{X}\mathscr{K}$ while 2m12m-1 is nondecreasing, which allow us to combine the characterizations (3.6) and (3.9) to deduce that the final term is nonpositive, yielding
@@
Here (XX~,VV~)(X-$\tilde${X},V-$\tilde${V}) corresponds to a uniform distribution centered at the center of mass for the system, and according to (7.7) this converges uniformly pointwise in $\Omega$ to the proposed steady state (7.8). From the fact that $\tilde{X},\tilde{V}\in T_{X}\mathscr{K}$, (3.6), (5.8), $\tilde{U}-\tilde{V}=U-V\in N_{X}\mathscr{K}$ and $\int_{0}^{1}\tilde{X}\mathop{}\!\mathrm{d}m=0=\int_{0}^{1}\tilde{U}\mathop{}\!\mathrm{d}m$ we derive the following relations,
@@
[…]